what is the value of "a" given the quadtratic equation:
2/a^2 + 3/a = -1
Thanks!

Answer by lwsshak3(11628)

what is the value of "a" given the quadtratic equation:
2/a^2 + 3/a = -1
**
let u=1/a
u^2=1/a^2
2u^2+3u=-1
2u^2+3u+1=0
(2u+1)(u+1)=0
u=-1/2
or
u=-1
..
u=-1/2=1/a
a=-2
or
u=-1=1/a
a=-1
..
a=-2 or -1